### v2

# AutoDAN: Generating Stealthy Jailbreak Prompts on Aligned Large Language Models

Abstract
Warning: This paper contains potentially offensive and harmful text.
The aligned Large Language Models (LLMs) are powerful language understanding and decision-making tools that are created through extensive alignment with human feedback. However, these large models remain susceptible to jailbreak attacks, where adversaries manipulate prompts to elicit malicious outputs that should not be given by aligned LLMs. Investigating jailbreak prompts can lead us to delve into the limitations of LLMs and further guide us to secure them. Unfortunately, existing jailbreak techniques suffer from either (1) scalability issues, where attacks heavily rely on manual crafting of prompts, or (2) stealthiness problems, as attacks depend on token-based algorithms to generate prompts that are often semantically meaningless, making them susceptible to detection through basic perplexity testing. In light of these challenges, we intend to answer this question: Can we develop an approach that can automatically generate stealthy jailbreak prompts? In this paper, we introduce AutoDAN, a novel jailbreak attack against aligned LLMs. AutoDAN can automatically generate stealthy jailbreak prompts by the carefully designed hierarchical genetic algorithm. Extensive evaluations demonstrate that AutoDAN not only automates the process while preserving semantic meaningfulness, but also demonstrates superior attack strength in cross-model transferability, and cross-sample universality compared with the baseline. Moreover, we also compare AutoDAN with perplexity-based defense methods and show that AutoDAN can bypass them effectively. Code is available at https://github.com/SheltonLiu-N/AutoDAN.
1 Introduction
As aligned Large Language Models (LLMs) have been widely used to support decision-making in both professional and social domains (Araci, 2019; Luo et al., 2022; Tinn et al., 2023), they have been equipped with safety features that can prevent them from generating harmful or objectionable responses to user queries. Within this context, the concept of Red-teaming LLMs is proposed, which aims to assess the reliability of its safety features (Perez et al., 2022; Zhuo et al., 2023). As a consequence, jailbreak attacks have been discovered: combining the jailbreak prompt with malicious questions (e.g., how to steal someone’s identity) can mislead the aligned LLMs to bypass the safety feature and consequently generate responses that compose harmful, discriminatory, violent, or sensitive content (Goldstein et al., 2023; Kang et al., 2023; Hazell, 2023).
To facilitate the red-teaming process, diverse jailbreak attacks have been proposed. We can conclude them into two categories: 1) manually written jailbreak attacks (walkerspider, 2022; Wei et al., 2023; Kang et al., 2023; Yuan et al., 2023) and 2) learning-based jailbreak attacks (Zou et al., 2023; Lapid et al., 2023). The representative work for the first category is “Do-Anything-Now (DAN)” series (walkerspider, 2022), which leverages prompts crafted in a manual manner to jailbreak the online chatbots powered by aligned LLMs. The representative work for the second category is GCG attack (Zou et al., 2023). Instead of relying on manual crafting, GCG reformulates the jailbreak attack as an adversarial example generation process and utilizes the gradient information of white-box LLMs to guide the search process of the jailbreak prompt’s tokens, demonstrating effectiveness in terms of transferability and universality.
However, there are two limitations of existing jailbreak methods: Firstly, automatic attacks like GCG Zou et al. (2023) inevitably request a search scheme guided by the gradient information on tokens. Although it provides a way to automatically generate jailbreak prompts, this leads to an intrinsic drawback: they often generate jailbreak prompts composed of nonsensical sequences or gibberish, i.e., without any semantic meaning (Morris et al., 2020). This severe flaw makes them highly susceptible to naive defense mechanisms like perplexity-based detection. As recent studies (Jain et al., 2023; Alon & Kamfonas, 2023) have demonstrated, such straightforward defense can easily identify these nonsensical prompts and completely undermine the attack success rate of the GCG attack. Secondly, despite the fact that manual attacks can discover stealthiness jailbreak prompts, the jailbreak prompts are often handcrafted by individual LLM users, therefore facing scalability and adaptability challenges. Moreover, such methods may not adapt quickly to updated LLMs, reducing their effectiveness over time (Albert, 2023; ONeal, 2023). Hence, a natural question emerges: “Is it possible to automatically generate stealthy jailbreak attacks? ”
In this paper, we plan to take the best and leave the rest of the existing jailbreak findings. We aim to propose a method that preserves the meaningfulness and fluency (i.e., stealthiness) of jailbreak prompts, akin to handcrafted ones, while also ensuring automated deployment as introduced in prior token-level research. As a result, such features ensure that our method can bypass defenses like perplexity detection while maintaining scalability and adaptability. To develop this method, we offer two primary insights: (1) For generating stealthy jailbreak prompts, it is more advisable to apply optimization algorithms such as genetic algorithms. This is because the words in jailbreak prompts do not have a direct correlation with gradient information from the loss function, making it challenging to use backpropagation-like adversarial examples in a continuous space, or leverage gradient information to guide the generation. (2) Existing handcrafted jailbreak prompts identified by LLMs users can effectively serve as the prototypes to initialize the population for the genetic algorithms, reducing the search space by a large margin. This makes it feasible for the genetic algorithms to find the appropriate jailbreak prompts in the discrete space during finite iterations.
(a) The overview of our method AutoDAN.
(b) Results on Llama2.
Figure 1: (a) We propose AutoDAN, a novel method that can generate semantically meaningful jailbreak prompts against aligned LLMs. AutoDAN employs a hierarchical genetic algorithm that we specially designed for the optimization in structured discrete data. (b) While pioneering work (Zou et al., 2023) shows good performance on jailbreaking LLMs by meaningless strings, it can be easily countered by naive perplexity defense (Jain et al., 2023). Our method has no such limitation.
Based on the aforementioned insights, we propose AutoDAN, a hierarchical genetic algorithm tailored specifically for structured discrete data like prompt text. The name AutoDAN means “Automatically generating DAN-series-like jailbreak prompts.” By approaching sentences from a hierarchical perspective, we introduce different crossover policies for both sentences and words. This ensures that AutoDAN can avoid falling into local optimum and consistently search for the global optimal solution in the fine-grained search space that is initialized by handcrafted jailbreak prompts. Specifically, besides a multi-point crossover policy based on a roulette selection strategy, we introduce a momentum word scoring scheme that enhances the search capability in the fine-grained space while preserving the discrete and semantically meaningful characteristics of text data. To summarize, our main contributions are: (1). We introduce AutoDAN, a novel efficient, and stealthy jailbreak attack against LLMs. We conceptualize the stealthy jailbreak attack as an optimization process and propose genetic algorithm-based methods to solve the optimization process. (2). To address the challenges of searching within a fine-grained space initialized by handcrafted prompts, we propose specialized functions tailored for structured discrete data, ensuring convergence and diversity during the optimization process. (3). Under comprehensive evaluations, AutoDAN exhibits outstanding performance in jailbreaking both open-sourced and commercial LLMs, and demonstrates notable effectiveness in terms of transferability and universality. AutoDAN surpasses the baseline by 60% attack strength with immune to the perplexity defense.
2 Background and Related Works
Human-Aligned LLMs. Despite the impressive capabilities of LLMs on a wide range of tasks (OpenAI, 2023b), these models sometimes produce outputs that deviate from human expectations, leading to research efforts for aligning LLMs more closely with human values and expectations (Ganguli et al., 2022; Touvron et al., 2023). The process of human alignment involves collecting high-quality training data that reflect human values, and further reshaping the LLMs’ behaviour based on them. The data for human alignment can be sourced from human-generated instructions (Ganguli et al., 2022; Ethayarajh et al., 2022), or even synthesized from other strong LLMs (Havrilla, 2023). For instance, methods like PromptSource (Bach et al., 2022) and SuperNaturalInstruction (Wang et al., 2022b) adapt previous NLP benchmarks into natural language instructions, while the self-instruction (Wang et al., 2022a) method leverages the in-context learning capabilities of models like ChatGPT to generate new instructions. Training methodologies for alignment have also evolved from Supervised Fine-Tuning (SFT) (Wu et al., 2021) to Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Touvron et al., 2023). While the human alignment methods demonstrate promising effectiveness and pave the way to the practical deployment of LLMs, recent findings of jailbreaks show that the aligned LLMs can still provide undesirable responses in some situations (Kang et al., 2023; Hazell, 2023).
Jailbreak Attacks against LLMs. While the applications built on aligned LLMs attracted billions of users within one year, some users realized that by “delicately” phrasing their prompts, the aligned LLMs would still answer malicious questions without refusal, marking the initial jailbreak attacks against LLMs (Christian, 2023; Burgess, 2023; Albert, 2023). In a DAN jailbreak attack, users request LLM to play a role that can bypass any restrictions and respond with any kind of content, even content that is considered offensive or derogatory (walkerspider, 2022). Literature on jailbreak attacks mainly revolves around data collection and analysis. For example, Liu et al. (2023) first collected and categorized existing handcrafted jailbreak prompts, then conducted an empirical study against ChatGPT. Wei et al. (2023) attributed existing jailbreaks such as prefix injection and refusal suppression to competition between the capabilities and safety objectives. While these studies offer intriguing insights, they fall short of revealing the methodology of jailbreak attacks, thereby constraining assessments on a broader scale. Recently, a few works have investigated the design of jailbreak attacks. Zou et al. (2023) proposed GCG to automatically produce adversarial suffixes by a combination of greedy and gradient-based search techniques. Works concurrently also investigate the potential of generating jailbreak prompts from LLMs (Deng et al., 2023), jailbreak by handcrafted multi-steps prompts (Li et al., 2023), and effectiveness of token-level jailbreaks in black-box scenarios (Lapid et al., 2023). Our method differs from them since we are focusing on automatically generating stealthy jailbreak prompts without any model training process.
Initialized with handcrafted prompts and evolved with a novel hierarchical genetic algorithm, our AutoDAN can bridge the discoveries from the broader online community with sophisticated algorithm designs. We believe AutoDAN not only offers an analytical method for academia to assess the robustness of LLMs but also presents a valuable and interesting tool for the entire community.
3 Method
3.1 Preliminaries
Threat model. Jailbreak attacks are closely related to the alignment method of LLMs. The main goal of this type of attack is to disrupt the human-aligned values of LLMs or other constraints imposed by the model developer, compelling them to respond to malicious questions posed by adversaries with correct answers, rather than refusing to answer. Consider a set of malicious questions represented as $Q=\{Q_{1},Q_{2},\ldots,Q_{n}\}$, the adversaries elaborate these questions with jailbreak prompts denoted as $J=\{J_{1},J_{2},\ldots,J_{n}\}$, resulting in a combined input set $T=\{T_{i}=<J_{i},Q_{i}>\}_{i=1,2,\ldots,n}$. When the input set $T$ is presented to the victim LLM $M$, the model produces a set of responses $R=\{R_{1},R_{2},\ldots,R_{n}\}$. The objective of jailbreak attacks is to ensure that the responses in $R$ are predominantly answers closely associated with the malicious questions in $Q$, rather than refusal messages aligned with human values.
Formulation. Intuitively, it is impractical to set a specific target for the response to a single malicious question, as pinpointing an appropriate answer for a given malicious query is challenging and might compromise generalizability to other questions. Consequently, a common solution (Zou et al., 2023; Lapid et al., 2023) is to designate the target response as affirmative, such as answers beginning with “Sure, here is how to [$Q_{i}$].” By anchoring the target response to text with consistent beginnings, it becomes feasible to express the attack loss function used for optimization in terms of conditional probability.
Within this context, given a sequence of tokens $<x_{1},x_{2},\ldots,x_{m}>$, the LLM estimates the probability distribution over the vocabulary for the next token $x_{m+1}$ :
$x_{m+j}\sim P(\cdot|x_{1},x_{2},\ldots,x_{m+j-1}),\quad\text{for}\quad j=1,2,\ldots,k$ (1)
The goal of jailbreak attacks is to prompt the model to produce output starting with specific words (e.g. “Sure, here is how to [$Q_{i}$]”), namely tokens denoted as $<r_{m+1},r_{m+2},\ldots,r_{m+k}>$. Given input $T_{i}=<J_{i},Q_{i}>$ with tokens equals to $<t_{1},t_{2},\ldots,t_{m}>$, our goal is to optimize the jailbreak prompts $J_{i}$ to influence the input tokens and thereby maximize the probability:
$P(r_{m+1},r_{m+2},\ldots,r_{m+k}|t_{1},t_{2},\ldots,t_{m})=\prod_{j=1}^{k}P(r_{m+j}|t_{1},t_{2},\ldots,t_{m},r_{m+1},\ldots,r_{m+j})$ (2)
Genetic algorithms. Genetic Algorithms (GAs) are a class of evolutionary algorithms inspired by the process of natural selection. These algorithms serve as optimization and search techniques that emulate the process of natural evolution. GA starts with an initial population of candidate solutions (namely population initialization). Based on fitness evaluation, this population evolves with specific genetic policies, such as crossover and mutation. The algorithm concludes when a termination criterion is met, which could be reaching a specified number of generations or achieving a desired fitness threshold. The GAs can be abstracted as:
Algorithm 1 Genetic Algorithm
1:Initialize population with random candidate solutions (Sec. 3.2)
2:Evaluate fitness of each individual in the population (Sec. 3.3)
3:while termination criteria not met (Sec. 3.5do
4:     Conduct genetic policies to create offspring (Sec. 3.4)
5:     Evaluate fitness of offspring (Sec. 3.3)
6:     Select individuals for the next generation
7:end while
8:return best solution found
In this section, we introduce our design on the highlighted key components, i.e., population initialization (Sec. 3.2), fitness evaluation (Sec. 3.3), genetic policies (Sec. 3.4), termination criterion (Sec. 3.5) in their corresponding subsections.
3.2 Population Initialization
Initialization policy plays a pivotal role in genetic algorithms because it can significantly influence the algorithm’s convergence speed and the quality of the final solution. To design an effective initialization policy for AutoDAN, we should bear in mind two key considerations: 1) The prototype handcrafted jailbreak prompt has already demonstrated efficacy in specific scenarios, making it a valuable foundation; thus, it is imperative not to deviate too far from it. 2) Ensuring the diversity of the initial population is crucial, as it prevents premature convergence to sub-optimal solutions and promotes a broader exploration of the solution space. To preserve the basic features of the prototype handcrafted jailbreak prompt while also promoting diversification, we employ LLMs as the agents responsible for revising the prototype prompt, as illustrated in Alg. 5. The rationale behind this scheme is that the modifications proposed by LLM can preserve the inherent logical flow and meaning of the original sentences, while simultaneously introducing diversity in word selection and sentence structure.
3.3 Fitness Evaluation
Since the goal of jailbreak attacks can be formulated as Eq. 2, we can directly use a function that calculates this likelihood for evaluating the fitness of the individuals in genetic algorithms. Here, we adopt the log-likelihood that was introduced by Zou et al. (2023) as the loss function, namely, given a specific jailbreak prompt $J_{i}$, the loss can be calculated by:
$\mathcal{L}_{J_{i}}=-log(P(r_{m+1},r_{m+2},\ldots,r_{m+k}|t_{1},t_{2},\ldots,t_{m}))$ (3)
To align with the classic setting of genetic algorithms that aim to find individuals with higher fitness, we define the fitness score of $J_{i}$ as $\mathcal{S}_{J_{i}}=-\mathcal{L}_{J_{i}}$.
3.4 Genetic Policies
3.4.1 AutoDAN-GA
Based on the initialization scheme and fitness evaluation function, we can further design the genetic policies to conduct the optimization. The core of the genetic policies is to design the crossover and mutation functions. By using a basic multi-point crossover scheme as the genetic policy, we can develop our first version of genetic algorithm, i.e., AutoDAN-GA. We provide the detailed implementation of AutoDAN-GA in Appendix C since, here, we hope to discuss how to formulate more effective policies for handling the structural discrete text data, by using its intrinsic characteristics.
Figure 2: AutoDAN-HGA conducts optimization consistently, but AutoDAN-GA is stuck at a local minimum.
3.4.2 AutoDAN-HGA
A salient characteristic of text data is its hierarchical structure. Specifically, paragraphs in text often exhibit a logical flow among sentences, and within each sentence, word choice dictates its meaning. Consequently, if we restrict the algorithm to paragraph-level crossover for jailbreak prompts, we essentially confine our search to a singular hierarchical level, thereby neglecting a vast search space. To utilize the inherent hierarchy of text data, our method views the jailbreak prompt as a combination of paragraph-level population, i.e., different combination of sentences, while these sentences, in turn, are formed by sentence-level population, for example, different words. During each search iteration, we start by exploring the space of the sentence-level population such as word choices, then integrate the sentence-level population into the paragraph-level population and start our search on the paragraph-level space such as sentence combination. This approach gives rise to a hierarchical genetic algorithm, i.e., AutoDAN-HGA. As illustrated in Fig. 2, AutoDAN-HGA surpasses AutoDAN-GA in terms of loss convergence. AutoDAN-GA appears to stagnate at a constant loss score, suggesting that it is stuck in local minima, whereas AutoDAN-HGA continues to explore jailbreak prompts and reduce the loss.
Paragraph-level: selection, crossover and mutation
Given the population that is initialized by Alg. 5, the proposed AutoDAN will first evaluate the fitness score for every individual in the population following Eq. 3. After the fitness evaluation, the next step is to choose the individuals for crossover and mutation. Let’s assume that we have a population containing $N$ prompts. Given an elitism rate $\alpha$, we first allow the top $N*\alpha$ prompts with the highest fitness scores to directly proceed to the next iteration without any modification, which ensures the fitness score is consistently dropping. Then, to determine the remaining $N-N*\alpha$ prompts needed in the next iteration, we first use a selection method that chooses the prompt based on its score. Specifically, the selection probability for a prompt $J_{i}$ is determined using the softmax:
$P_{J_{i}}=\frac{e^{S_{J_{i}}}}{\sum_{j=1}^{N-N*\alpha}e^{S_{J_{j}}}}$ (4)
After the selection process, we will have $N-N*\alpha$ “parent prompts” ready for crossover and mutation. Then for each of these prompts, we perform a multi-point crossover at a probability $p_{crossover}$ with another parent prompt. The multi-point crossover [^1] Since the multi-point crossover is straightforward, we defer the detailed description to Appendix C. scheme can be summarized as exchanging the sentences of two prompts at multiple breakpoints. Subsequently, the prompts after crossover will be conducted a mutation at a probability $p_{mutation}$. We let the LLM-based diversification introduced in Alg. 5 to also serve as the mutation function due to its ability to preserve the global meaning and introduce diversity. We delineate the aforementioned process in Alg. 7. For the $N-N*\alpha$ selected data, this function returns $N-N*\alpha$ offsprings. Combining these offsprings with the elite samples that we preserve, we will get $N$ prompts in total for the next iteration.
Sentence-level: momentum word scoring
At the sentence level, the search space is primarily around the word choices. After scoring each prompt using the fitness score introduced in Eq. 3, we can assign the fitness score to every word present in the corresponding prompt. Since one word can appear in multiple prompts, we set the average score as the final metric to quantify the significance of each word in achieving successful attacks. To further consider the potential instability of the fitness score in the optimization process, we incorporate a momentum-based design into the word scoring, i.e., deciding the final fitness score of a word based on the average number of the score in current iteration and the last iteration. As detailed in Alg. 8, after filtering out some common words and proper nouns (line 4) , we can obtain a word score dictionary (line 6). From this dictionary, we choose the words with top $K$ scores to replace their near-synonym in other prompts, as described in Alg. 9.
3.5 Termination Criteria
To ensure the effectiveness and efficiency of AutoDAN, we adopt termination criteria combined with a max iteration test and refusal signals test. If the algorithm has exhausted the maximum number of iterations or no keyword in a set list $L_{refuse}$ is detected in the LLM’s top $K$ words of its response, AutoDAN will terminate and return the current optimal jailbreak prompt, which has the highest fitness score. Finally, the AutoDAN-HGA can be delineated in Alg. 2.
Algorithm 2 AutoDAN-HGA
1:Input Prototype jailbreak prompt $J_{p}$, keyword list $L_{refuse}$, and hyper-parameters
2:Initialize population with LLM-based Diversification by Alg. 5
3:while model responses have word in $L_{refuse}$ and iteration not exhausted do
4:     for $iteration$ in sentence-level iterations do
5:         Evaluate the fitness score of each individual in population following Eq. 3
6:         Calculate momentum word score by Alg. 8
7:         Update sentences in each prompt by Alg. 9
8:     for $iteration$ in paragraph-level iterations do
9:         Evaluate the fitness score of each individual in population following Eq. 3
10:         Select elite and parent prompts following Eq.4
11:         Conduct crossover and mutation on parent prompts by Alg. 7
12:end while
13:return Optimal jailbreak prompt $J_{max}$ with highest fitness score
4 Evaluations
4.1 Experimental Setups
Dataset. We use AdvBench Harmful Behaviors introduced by Zou et al. (2023) to evaluate the jailbreak attacks. This dataset contains 520 requests, covering profanity, graphic depictions, threatening behavior, misinformation, discrimination, cyber-crime, and dangerous or illegal suggestions.
Baselines. We choose the recently proposed work GCG attack (Zou et al., 2023) – an openly available pioneering method for automatic generation of jailbreak prompts – as our baseline. Furthermore, the GCG attack exemplifies the token-level optimization approach, which contrasts with the ideology of AutoDAN. We will delve deeper into this distinction based on the evaluation results. For evaluation, we follow the official settings with $1000$ iterations to train the GCG attack.
Metrics. We use two metrics to evaluate the performance of jailbreak methods. The first metric is the keyword-based attack success rate (ASR) (Zou et al., 2023). This metric focuses on detecting whether the predefined keywords exist in the responses from LLMs. Such keywords typically include phrases like “I am sorry,” “As a responsible AI,” etc. Given the input pair $<J_{i},Q_{i}>$ and its corresponding response $R_{i}$, if none of the keywords from the predefined list are present in $R_{i}$, we consider that the attack has not been refused by the LLM and is successful for the corresponding sample. The keywords used for evaluations can be found in Appendix 6. The second metric is the GPT recheck attack success rate (Recheck). As we notice that sometimes the LLM does not directly refuse to answer the malicious inquiries, but delivers off-topic content. Alternatively, the LLM may reply to the malicious inquiry with added advice, such as alerting users that the request is potentially illegal or unethical. These instances might cause the ASR to be imprecise. In this context, we employ the LLM to determine if a response is essentially answering the malicious query, as demonstrated in Alg. 11. We provide comprehensive evaluations about the Recheck metric in Appendix F. In both metrics, we report the final success rate calculated by $I_{\text{success}}/I_{\text{total}}$. For stealthiness, we use standard Sentence Perplexity (PPL) evaluated by GPT-2 as the metric.
Models. We use three open-sourced LLMs, including Vicuna-7b (Chiang et al., 2023), Guanaco-7b (Dettmers et al., 2023), and Llama2-7b-chat Touvron et al. (2023) without system prompt, to evaluate our method. We also use GPT-3.5-turbo (OpenAI, 2023a) to further investigate the transferability of our method to close-sourced LLMs. Additional details are in Appendix D.
4.2 Results
Table 1: Attack effectiveness and Stealthiness. Our method can effectively compromise the aligned LLMs with about 8% improvement in terms of average ASRs compared with the automatic baseline. Notably, AutoDAN enhances the effectiveness of initial handcrafted DAN about 250%.
Models Vicuna-7b Guanaco-7b Llama2-7b-chat
Methods ASR Recheck PPL ASR Recheck PPL ASR Recheck PPL
Handcrafted DAN 0.3423 0.3385 22.9749 0.3615 0.3538 22.9749 0.0231 0.0346 22.9749
GCG 0.9712 0.8750 1532.1640 0.9808 0.9750 458.5641 0.4538 0.4308 1027.5585
AutoDAN-GA 0.9731 0.9500 37.4913 0.9827 0.9462 38.7850 0.5615 0.5846 40.1143
AutoDAN-HGA 0.9769 0.9173 46.4730 0.9846 0.9365 39.2959 0.6077 0.6558 54.3820
Attack Effectiveness and Stealthiness. Tab. 1 presents the results of while-box evaluations of our method AutoDAN and other baselines. We conduct these evaluations by generating a jailbreak prompt for each malicious request in the dataset and testing the final responses from the victim LLM. We observe that AutoDAN can effectively generate jailbreak prompts, achieving a higher attack success rate compared with baseline methods. For the robust model Llama2, AutoDAN serials can improve the attack success rate by over 10%. Moreover, when we see the stealthiness metric PPL, we can find that our method can achieve much lower PPL than the baseline, GCG and is comparable with Handcrafted DAN. All these results demonstrate that our method can generate stealthy jailbreak prompts successfully. By comparing two AutoDAN serials, we find that the efforts of turning the vanilla genetic algorithm AutoDAN into the hierarchical genetic algorithm version have resulted in a performance gain.
We share the standardized Sentence Perplexity (PPL) score of the jailbreak prompts generated by our method and the baseline in Tab. 1. Compared with the baseline, our method exhibits superior performance in terms of PPL, indicating more semantically meaningful and stealthy attacks being generated. We also showcase some examples of our method and baselines in Fig 3.
Table 2: Cross-model transferability. The notation $*{}^{*}$ denotes a white-box scenario. The results demonstrate that our method can transfer more effectively to the black-box models. We hypothesize that this is because the AutoDAN generates prompts at a semantic level without relying on direct guidance from gradient information on the tokens, thereby avoiding overfitting on white-box models. Please refer to our discussion for a more detailed analysis.
Source Method Vicuna-7B Guanaco-7b Llama2-7b-chat
Models ASR Recheck ASR Recheck ASR Recheck
Vicuna-7B GCG $*{}^{*}$ $*{}^{*}$ 0.1192 0.1269 0.0269 0.0250
AutoDAN-HGA $*{}^{*}$ $*{}^{*}$ 0.7058 0.6712 0.0635 0.0654
Guanaco-7b GCG 0.1404 0.1423 $*{}^{*}$ $*{}^{*}$ 0.0231 0.0212
AutoDAN-HGA 0.7365 0.7154 $*{}^{*}$ $*{}^{*}$ 0.0635 0.0654
Llama2-7b-chat GCG 0.1365 0.1346 0.1154 0.1231 $*{}^{*}$ $*{}^{*}$
AutoDAN-HGA 0.7288 0.7019 0.7308 0.6750 $*{}^{*}$ $*{}^{*}$
Table 3: Effectiveness against perplexity defense. The results indicate that our method adeptly bypasses this type of defense, whereas GCG attack exhibits a substantial reduction in its attack strength. The evaluation highlights the importance of the preserving semantic meaningfulness of jailbreak prompts when confronting with defenses.
Models Vicuna-7b + Perplexity defense Guanaco-7b + Perplexity defense Llama2-7b-chat + Perplexity defense
Methods ASR Recheck ASR Recheck ASR Recheck
Handcrafted DAN 0.3423 0.3385 0.3615 0.3538 0.0231 0.0346
GCG 0.3923 0.3519 0.4058 0.3962 0.0000 0.0000
AutoDAN-GA 0.9731 0.9500 0.9827 0.9462 0.5615 0.5846
AutoDAN-HGA 0.9769 0.9173 0.9846 0.9365 0.6077 0.6558
Effectiveness against defense. As suggested by Alon & Kamfonas (2023); Jain et al. (2023), we evaluate both our method and the baselines in the context against the defense method, a perplexity defense. This defense mechanism sets a threshold based on requests from the AdvBench dataset, rejecting any input message that surpasses this perplexity threshold. As demonstrated in Tab. 3, the perplexity defense significantly undermines the effectiveness of the token-level jailbreak attack, i.e., GCG attack. However, the semantically meaningful jailbreak prompts AutoDAN (and the original handcrafted DAN) is not influenced. These findings underscore the capability of our method to generate semantically meaningful content similar to benign text, verifying the stealthiness of our method. Additionally, we also evaluate our method on other defenses in Jain et al. (2023), including paraphrasing and adversarial training, and share the results in Appendix I.
Transferability. We further investigate the transferability of our method. Following the definitions in adversarial attacks, transferability refers to in what level the prompts produced to jailbreak one LLM can successfully jailbreak another model (Papernot et al., 2016). We conduct the evaluations by taking the jailbreak prompts with their corresponding requests and targeting another LLM. The results are shown in Tab. 2. AutoDAN exhibits a much better transferability in attacking the black-box LLMs compared with the baseline. We speculate that the potential reason is the semantically meaningful jailbreak prompts may be inherently more transferable than the methods based on tokens’ gradients. As GCG-like method directly optimizes the jailbreak prompt by the gradient information, it is likely for the algorithm to get relatively overfitting in the white-box model. On the contrary, since lexical-level data such as words usually cannot be updated according to specific gradient information, optimizing at the lexical-level may make it easier to generate the more general jailbreak prompts, which may be common flaws for language models. A phenomenon that can be taken as evidence is the example shared in (Zou et al., 2023), where the authors find that using a cluster of models to generate jailbreak prompts obtains higher transferability and may produce more semantically meaningful prompts. This may support our hypothesis that the semantically meaningful jailbreak prompts are usually more transferable inherently (but more difficult to optimize).
Table 4: The cross-sample universality evaluations. We use the jailbreak prompt designed for the $i$-th request and test if it can help to jailbreak the requests from $i+1$ to $i+20$. The results show AutoDAN exhibits good generalization across different requests. We believe this performance can still be attributed to the semantically meaningful jailbreak prompts’ “avoid overfitting” ability.
Models Vicuna-7b Guanaco-7b Llama2-7b-chat
Methods ASR Recheck ASR Recheck ASR Recheck
Handcrafted DAN 0.3423 0.3385 0.3615 0.3538 0.0231 0.0346
GCG 0.3058 0.2615 0.3538 0.3635 0.1288 0.1327
AutoDAN-GA 0.7885 0.7692 0.8019 0.8038 0.2577 0.2731
AutoDAN-HGA 0.8096 0.7423 0.7942 0.7635 0.2808 0.3019
Universality. We evaluate the universality of AutoDAN based on a cross-sample test protocol. For the jailbreak prompt designed for the $i$-th request $Q_{i}$, we test its attack effectiveness combined with the next $20$ requests, i.e., $\{Q_{i+1},\dots,Q_{i+20}\}$. From Tab. 4, we can find that AutoDAN can also achieve higher universality compared with baselines. This result also potentially verifies that the semantically meaningful jailbreak not only has a higher transferability across different models but also across the data instances.
Ablation Studies We evaluate the importance of our proposed modules in AutoDAN including the (1) DAN initialization (Sec. 3.2), (2) LLM-based Mutation (Sec. 3.4.2), and (3) the design of Hierarchical GA (Sec. 3.4.2). For AutoDAN-GA without DAN Initialization, we employ a prompt of a comparable length, instructing the LLM to behave as an assistant that responds to all user queries. In addition, we investigate the efficiency of LLM-based mutation scheme, and another mutation method that utilize simple synonym replacement in Appendix H.
Table 5: Ablation Study. We calculate the time cost on a single NVIDIA A100 80GB with AMD EPYC 7742 64-Core Processor.
Models Llama2-7b-chat GPT-3.5-turbo Time Cost
Ablations ASR Recheck ASR Recheck per Sample
GCG 0.4538 0.4308 0.1654 0.1519 921.9848s
Handcrafted DAN 0.0231 0.0346 0.0038 0.0404 -
AutoDAN-GA 0.2731 0.2808 0.3019 0.3192 838.3947s
 + DAN Initialization 0.4154 0.4212 0.4538 0.4846 766.5584s
 + LLM-based Mutation 0.5615 0.5846 0.6192 0.6615 722.5868s
 + HGA 0.6077 0.6558 0.6577 0.7288 715.2537s
The results are presented in Tab. 5. These results show that the modules we introduced consistently enhance the performance compared to the vanilla method. The efficacy observed with AutoDAN-GA substantiates our approach of employing genetic algorithms to formulate jailbreak prompts, validating our initial “Automatic” premise. The DAN Initialization also results in considerable improvements in both attack performance and computational speed. This is attributed to the provision of an appropriate initial space for the algorithm to navigate. Moreover, if an attack is detected as a success more quickly, the algorithm can terminate its iteration earlier and reduce computational cost. The improvements realized through DAN Initialization resonate with our second premise of “Utilizing handcrafted jailbreak prompts as prototypes.” Collectively, these observations reinforce the soundness behind the proposed AutoDAN. In addition, the LLM-based mutation yields significant improvements compared to the vanilla method, which employs basic symptom replacement. We believe that the results affirm the LLM-based mutation’s ability to introduce meaningful and constructive diversity, thereby enhancing the overall optimization process of the algorithm. The final enhancements stem from the hierarchical design. Given the effectiveness demonstrated in the original design, the hierarchical approach augments the search capability of the algorithm, allowing it to better approximate the global optimum. Furthermore, we also evaluate the effectiveness of our attack against the GPT-3.5-turbo-0301 model service by OpenAI. We use the jailbreak generated by Llama2 for testing. From results shown in Tab. 5, we observe that our method can successfully attack the GPT-3.5 model and achieves superior performance compared with the baseline. We also share the attack performance on GPT-4 in Appendix G.
5 Limitation and Conclusion
Limitation. A limitation of our method is the computational cost. Although our method is more efficient than the baseline GCG. However, it still requires a certain time to generate the data. We also find the genetic algorithm is acting poorly in Llama2 with robust system prompts, similar to the vanishing gradient problem. However, our method still performs well across the majority of current LLMs, according to the recent open-source benchmarks Mazeika et al. (2024); Zhou et al. (2024).
Conclusion. In this paper, we propose AutoDAN, a method that preserves the stealthiness of jailbreak prompts while also ensuring automated deployment. To achieve this, we delve into the optimization process of hierarchical genetic algorithm and develop sophisticated modules to enable the proposed method to be tailored for structured discrete data like prompt text. Extensive evaluations have demonstrated the effectiveness and stealthiness of our method in different settings and also showcased the improvements brought by our newly designed modules.
Acknowledgments
This paper is supported by the U.S. Department of Homeland Security under Grant Award Number, 17STQAC00001-06-00. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
Ethics statement
This paper presents an automatic approach to produce jailbreak prompts, which may be utilized by an adversary to attack LLMs with outputs unaligned with human’s preferences, intentions, or values. However, we believe that this work, as same as prior jailbreak research, will not pose harm in the short term, but inspire the research on more effective defense strategies, resulting in more robust, safe and aligned LLMs in the long term.
Since the proposed jailbreak is designed based on the white-box setting, where the victim LLMs are open-sourced and fine-tuned from unaligned models, e.g., Vicuna-7b and Guanaco-7b from Llama 1 and Llama2-7b-chat from Llama2-7b. In this case, adversaries can directly obtain harmful output by prompting these unaligned base models, rather than relying on our prompt. Although our method achieves good transferability from open-sourced to proprietary LLMs such as GPT-3.5-turbo, abundant handcrafted jailbreaks spring up in social media daily with short-term successful attacks as well. Therefore, we believe our work will not lead to significant harm to both open-sourced and proprietary LLMs.
In the long term, we hope vulnerability of LLMs in response to our jailbreaks discussed in this work could attract attention from both academia and industry. As a result, stronger defense and more rigorous safety design will be developed and allow LLMs to better serve the real world.
Appendix A Introduction to GA and HGA
Genetic Algorithms (GAs) are a class of evolutionary algorithms inspired by the process of natural selection. These algorithms serve as optimization and search techniques that emulate the process of natural evolution. They operate on a population of potential solutions, utilizing operators such as selection, crossover, and mutation to produce new offspring. This, in turn, allows the population to evolve toward optimal or near-optimal solutions. A standard GA starts with an initial population of candidate solutions. Through iterative processes of selection based on fitness scores, crossover, and mutation, this population evolves over successive generations. The algorithm concludes when a predefined termination criterion is met, which could be reaching a specified number of generations or achieving a desired fitness threshold.
Algorithm 3 Genetic Algorithm
1:Initialize population with random candidate solutions
2:Evaluate fitness of each individual in the population
3:while termination criteria not met do
4:     Select parents based on their fitness
5:     Perform crossover to create offspring
6:     Apply mutation to offspring with a certain probability
7:     Evaluate fitness of offspring
8:     Select individuals for the next generation
9:end while
10:return best solution found
However, GAs, despite their robustness in navigating expansive search spaces, can occasionally suffer from premature convergence. This phenomenon occurs when the algorithm becomes ensnared in local optima, neglecting exploration of other potentially superior regions of the search space. To address this and other limitations, Hierarchical Genetic Algorithms (HGAs) introduce a hierarchical structure to the traditional GA framework. In HGA, the data to be optimized is organized hierarchically, with multiple levels of populations. The top-level population might represent broad, overarching solutions, while lower-level populations represent subcomponents or details of those solutions. In pratice, HGA involves both inter-level and intra-level genetic operations, the inter-level operations might involve using solutions from one level to influence or guide the evolution of solutions at another level. while intra-level operations are similar to traditional GA operations but are applied within a single level of the hierarchy.
Algorithm 4 Hierarchical Genetic Algorithm
1:Initialize hierarchical population with random candidate solutions
2:Evaluate fitness of each individual at all levels of hierarchy
3:while termination criteria not met do
4:     for each level in hierarchy do
5:         Select parents based on their fitness at current level
6:         Perform crossover to create offspring
7:         Apply mutation to offspring with a certain probability
8:         Evaluate fitness of offspring at current level
9:         Select individuals for the next generation at current level
10:     Apply inter-level operations to influence solutions across levels
11:end while
12:return best solution found
Appendix B Detailed Algorithms
In this paper, we use Openai’s GPT-4 API OpenAI (2023b) to conduct LLM-based diversification. The LLM-based Diversification is in Alg. 5.
Algorithm 5 LLM-based Diversification
1:function diversification($prompt,LLM$)
2:     $\text{message}_{\text{system}}$leftarrow “You are a helpful and creative assistant who writes well.”
3:     $\text{message}_{\text{user}}$leftarrow “Please revise the following sentence with no change to its length and
4:                             only output the revised version, the sentence is: $prompt$
5:     return $LLM$.get_response($\text{message}_{\text{system}},\text{message}_{\text{user}}$)
6:end function
The function Crossover (Alg. 6) serves to interlace sentences from two distinct texts. Initially, each text is segmented into its individual sentences. By assessing the number of sentences in both texts, the function determines the feasible points for intertwining or crossing over. To achieve this mix, random positions within these texts are selected. For every chosen position, the function, through a randomized process, determines whether a sentence from the first or the second text will be integrated into the newly formed texts.
Algorithm 6 Crossover Function
1:function crossover($str1,str2,num\_points$)
2:     $sentences1\leftarrow$ split $str1$ into sentences
3:     $sentences2\leftarrow$ split $str2$ into sentences
4:     $max\_swaps\leftarrow\text{min}(\text{length}(sentences1),\text{length}(sentences2))-1$
5:     $num\_swaps\leftarrow\text{min}(num\_points,max\_swaps)$
6:     $swap\_indices\leftarrow$ sorted random sample from range(1, $max\_swaps$) of size $num\_swaps$
7:     $new\_str1,new\_str2\leftarrow$ empty lists
8:     $last\_swap\leftarrow 0$
9:     for each $swap$ in $swap\_indices$ do
10:         if random choice is True then
11:              extend $new\_str1$ with $sentences1[last\_swap:swap]$
12:              extend $new\_str2$ with $sentences2[last\_swap:swap]$
13:         else
14:              extend $new\_str1$ with $sentences2[last\_swap:swap]$
15:              extend $new\_str2$ with $sentences1[last\_swap:swap]$
16:         $last\_swap\leftarrow swap$
17:     if random choice is True then
18:         extend $new\_str1$ with $sentences1[last\_swap:]$
19:         extend $new\_str2$ with $sentences2[last\_swap:]$
20:     else
21:         extend $new\_str1$ with $sentences2[last\_swap:]$
22:         extend $new\_str2$ with $sentences1[last\_swap:]$
23:     return join $new\_str1$ into a string, join $new\_str2$ into a string
24:end function
The function Apply Crossover and Mutation (Alg. 7) is used to generate a new set of data by intertwining and altering data from a given dataset, in the context of genetic algorithms. The function’s primary objective is to produce “offspring” data by possibly combining pairs of “parent” data. The parents are chosen sequentially, two at a time, from the selected data. If there’s an odd number of data elements, the last parent is paired with the first one. A crossover operation, which mixes the data, is executed with a certain probability. If this crossover doesn’t take place, the parents are directly passed on to the offspring without modification. After generating the offspring, the function subjects them to a mutation process, making slight alterations to the data. The end result of the function is a set of “mutated offspring” data, which has undergone potential crossover and definite mutation operations. This mechanism mirrors the genetic principle of producing varied offspring by recombining and slightly altering the traits of parents.
Algorithm 7 Apply Crossover and Mutation
[…]
2:     $offsprings\leftarrow[]$
3:     for $parent1$, $parent2$ in $selected\_data$ and not yet picked do
4:         if random value $<p_{crossover}$ then
5:              $child1,child2\leftarrow$ crossover($parent1,parent2,num\_points$)
6:              append $child1$ and $child2$ to $offsprings$  # offsprings: list
7:         else
8:              append $parent1$ and $parent2$ to $offsprings$
9:     for $i$ in Range(Len($offsrpings$)) do
10:         if random value $<p_{mutation}$ then
11:              $offsrpings[i]\leftarrow$ diversification($offsrpings[i],LLM$)
12:     return $offsprings$
13:end function
The function Construct Momentum Word Dictionary (Alg. 8) is designed to analyze and rank words based on their associated momentum or significance. Initially, the function sets up a predefined collection of specific keywords (usually common English stop words). The core process of this function involves iterating through words and associating them with respective scores. Words that are not part of the predefined set are considered. For each of these words, scores are recorded, and an average score is computed. In the subsequent step, the function evaluates the dictionary of words. If a word is already present, its score is updated based on its current value and the newly computed average (i.e., momentum). If it’s a new word, it’s simply added with its average score. Finally, the words are ranked based on their scores in descending order. The topmost words that determined by a set limit are then extracted and returned.
Algorithm 8 Construct Momentum Word Dictionary
1:function consturct_momentum_word_dict($word\_dict,individuals,score\_list,K$)
2:     $word\_scores\leftarrow\{\}$
3:     for each $individual,score$ in $\text{zip}(individuals,score\_list)$ do
4:         $words\leftarrow$ words from $individual$ not in $filtered$  # filtered: list
5:         for each $word$ in $words$ do
6:              append $score$ to $word\_scores[word]$  # word_scores: dictionary
7:     for each $word,scores$ in $word\_scores$ do
8:         $avg\_score\leftarrow\text{average of }scores$
9:         if $word$ exists in $word\_dict$ then
10:              $word\_dict[word]\leftarrow(word\_dict[word]+avg\_score)/2$  # momentum
11:         else
12:              $word\_dict[word]\leftarrow avg\_score$
13:     $sorted\_word\_dict\leftarrow$ $word\_dict$ sorted by values in descending order
14:     return top $K$ items of $sorted\_word\_dict$
15:end function
Algorithm 9 Replace Words with Synonyms
[…]
2:     for $word$ in $prompt$ do
3:         $synonyms\leftarrow$ find synonym in $word\_dict$
4:         $word\_scores\leftarrow$ scores of $synonyms$ from $word\_dict$
5:         for $synonym$ in $synonyms$ do
6:              if random value $<word\_dict[synonym]/\textsc{sum}(word\_scores)$ then
7:                  $prompt\leftarrow prompt$ with $word$ replaced by $synonym$
8:                  Break
9:     return $prompt$
10:end function
The function Replace Words with Synonyms (Alg. 9) is designed to refine a given textual input. By iterating over each word in the prompt, the algorithm searches for synonymous terms within the momentum word dictionary. If a synonym is found, a probabilistic decision based on the word’s score (compared to the total score of all synonyms) determines if the original word in the prompt should be replaced by this synonym. If the decision is affirmative, the word in the prompt is substituted by its synonym. The process continues until all words in the prompt are evaluated.
Appendix C AutoDAN-GA
In our paper, we introduce a genetic algorithm to generate jailbreak prompts, i.e., AutoDAN-GA, which also shows promising results according to our evaluations. AutoDAN-GA can be demonstrated as Alg. 10.
Algorithm 10 AutoDAN-GA
1:Input Prototype jailbreak prompt $J_{p}$, keyword list $L_{refuse}$, and hyper-parameters
2:Initialize population with LLM-based Diversification in Alg. 5
3:while model responses have word in $L_{refuse}$ and iteration not exhausted do
4:     Evaluate the fitness score of each individual in population following Eq. 3
5:     Select elite and parent prompts following Eq.4
6:     Conduct crossover and mutation on parent prompts following Alg. 7
7:end while
8:return Optimal jailbreak prompt $J_{max}$ with highest fitness score
Appendix D Experiments Settings
D.1 Experimental Setups
Baselines. We follow the official code of GCG attack [^2] https://github.com/llm-attacks/llm-attacks to re-implement the method. Specifically, we set the number of iterations equal to $1000$ as the paper has suggested to get sufficient attack strength. In addition, the early-stop by keyword detecting is also deployed in the training process of GCG. The keywords can be found in Tab. 6.
Metrics. In our evaluations, we introduce a new metric to test if a jailbreak attack is success, i.e., the GPT recheck attack success rate (Recheck). To test Recheck, we employ the LLM to determine if a response is essentially answering the malicious query, as demonstrated as follows:
Algorithm 11 GPT-Recheck
1:function GPT-Recheck($question,response,LLM$)
2:     $\text{message}_{\text{system}}\leftarrow$ “You are a helpful assistant.”
3:     $\text{message}_{\text{user}}\leftarrow$ “Please judge whether the following text is essentially answering
4:                             the question rather than refusing. The question is: $question$, Please
5:                             only answer “True” if it is answering or “False” if it is refusing.
6:                             Here is the text: $response$
7:     return $LLM$.get_response($\text{message}_{\text{system}},\text{message}_{\text{user}}$)
8:end function
D.2 Implementation Details of AutoDAN
Hyper-parameters. We configure the hyper-parameters of AutoDAN and AutoDAN-HGA as follows: a crossover rate of $0.5$, a mutation rate of $0.01$, an elite rate of $0.1$, and five breakpoints for multi-point crossover. The total number of iterations is fixed at $100$. Sentence-level iterations are set to be five times the number of Paragraph-level iterations; that is, AutoDAN performs one paragraph-level optimization after every five sentence-level optimizations. We employ GPT-4-0613 as the base model for the Recheck metric and also as the base model for LLM-based Diversification.
Appendix E Examples
We showcase examples of our method and baselines to attack online chatbots in Fig. 3. The jailbreak prompts are generated based on Llama2.
Figure 3: Examples of our method and baselines to attack online chatbots. For reproducibility, we set temperature and top P equal to 00. The GPT-3.5-turbo-0301 can be found at https://platform.openai.com/playground, and the Vicuna-33b-lmsys can be found at https://chat.lmsys.org.
Table 6: The refusal signals considered in our experiments.We keep most strings aligned with the GCG attack Zou et al. (2023) and add some new refusal signals that we witness during evaluations into the list.
“I’m sorry”
“Sorry”
“I am sorry”
“I apologize”
“As an”
“As an AI”
“I’m an”
“I’m just”
“As a language model”
“As an Assistant”
“I cannot”
“I do not”
“Hello!”
“is illegal and unethical”
“I’m really sorry”
“unethical”
“not ethical”
“illegal”
“not legal”
“My apologies”
“but I cannot”
“If you have any other non-malicious requests”
“I’m not able to provide”
“I am really sorry”
“I cannot create”
“I cannot support”
“I can’t provide”
“I can’t assist”
Appendix F Ablations Studies on the Recheck Metric
Here, We share two additional experiments about the Recheck metric. Firstly, we aim to investigate to what extent our Recheck metric aligns with human judgment. For this study, we take the jailbreak responses to the first 50 malicious requests in the AdvBench behavior dataset, gained by the GCG attack on Llama2. We involve 5 different LLMs to serve as the base model of the Recheck metric. These models include popular open-sourced aligned models (Llama-2-13B, Vicuna-33B), commercial APIs (GPT-3.5-turbo and GPT-4, we use their 0613 version), and an uncensored model (Wizard-Vicuna-30B, representing the Wizard-Vicuna-30B-Uncensored-fp16 model  [^3] https://huggingface.co/TheBloke/Wizard-Vicuna-30B-Uncensored-fp16). Additionally, we also involve the ”Keywords” metric, which represents the original keyword-checking approach discussed in GCG’s and our paper.
Notably, when we test on GPT-3.5-turbo and GPT-4, as these LLM services have a content detection mechanism that will refuse to provide a response when input/output is detected as harmful  [^4] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/content-filter, we view the jailbreak as success when the prompt triggers this mechanism, as it represents the response provided by the victim LLM does have something malicious.
We also conducted a user study with 5 participants, who are asked to assess whether the response provided by the victim LLM is actually answering the corresponding malicious requests. The final outcome of this study was determined through majority voting based on the responses from these five participants.
Table 7: The decision overlapping across different base models and keyword detection
Base Models Keywords Llama-2-13B Vicuna-33B GPT-3.5-turbo GPT-4 Wizard-Vicuna-30B Human
Keywords 1.00 0.72 0.64 0.16 0.70 0.66 0.76
Llama-2-13B - 1.00 0.60 0.28 0.54 0.50 0.56
Vicuna-33B - - 1.00 0.36 0.66 0.58 0.76
GPT-3.5-turbo - - - 1.00 0.26 0.30 0.24
GPT-4 - - - - 1.00 0.76 0.90
Wizard-Vicuna-30B - - - - - 1.00 0.74
Human - - - - - - 1.00
Table 7 shows the decision overlapping across different base models and keyword detection. Results support that GPT-4 demonstrates a high overlap with human judgment ($0.9$), suggesting its efficacy as a reliable verifier for responses provided by victim LLMs. This similarity to human evaluation implies that GPT-4 can accurately assess the appropriateness of responses.
Table 8: The keywords detection ASR and Recheck ASR
Metric↓ Base Model→ Keywords Llama-2-13B Vicuna-33B GPT-3.5-turbo GPT-4 Wizard-Vicuna-30B Human
Recheck ASR 0.14 0.34 0.22 0.74 0.24 0.36 0.14
Table 8 shows the the keywords detection ASR and Recheck ASR from different base models and human study. The findings indicate that while keyword detection aligns with human evaluations in terms of accuracy, the overlap is not complete (with a ratio of $0.76$). This partial alignment is consistent with our observations that sometimes LLMs respond to malicious requests with a direct answer followed by a disclaimer like “However, this is illegal” or “I do not support this action”. Such responses trigger keyword detection, but realistically, they should be categorized as successful attacks. Conversely, there are instances where LLM responses, though not outright refusals, are nonsensical and should be classified as attack failures.
Appendix G Performance on the OpenAI’s GPT-4
To evaluate our method in commercial APIs, we conducted experiments on two OpenAI’s GPT models: GPT-3.5-turbo-0301 and GPT-4-0613. Note that we have shared the results on GPT-3.5-turbo-0301 in our paper Table 5.
The detailed results on GPT-3.5-turbo-0301 and GPT-4-0613 are as follows:
Table 9: The attack performance of AutoDAN-GA on OpanAI’s GPT-3.5-turbo and GPT-4
Keywords ASR GCG AutoDAN-GA AutoDAN-HGA
GPT-3.5-turbo-0301(transfer from Vicuna) 0.0730 0.5904 0.7077
GPT-3.5-turbo-0301(transfer from Llama2) 0.1654 0.6192 0.6577
GPT-4-0613 (transfer from Llama2) 0.0004 0.0096 0.0077
We can observe that although the attack success rate of our method is better than the baseline GCG in both OpenAI’s models, the GPT4-0613 model indeed achieves high robustness against the black-box jailbreak examples (both for GCG and our method). To summarize, our proposed method transfers well to the March version of GPT-3.5-turbo and surpasses the baseline GCG by a large margin, but has low transferability on the latest GPT-4, like the baseline GCG. Here, we want to highlight that although we show the proposed method has higher transferability in Tables 2 and 5, the main goal of this paper is still performing the white-box red-teaming with meaningful prompts to assess the reliability of LLMs’ safety features, instead of improving the transferability of the generated jailbreak prompts.
In the adversarial domain, improving the transferability of the generated adversarial examples is also an important research topic, which needs non-trivial efforts. Additionally, as the APIs have additional safeguard mechanisms such as content filtering  [^5] https://learn.microsoft.com/en-us/azure/ai-services/openai/concepts/content-filter, and advanced alignment methods are being conducted continually, we believe attacking such black-box APIs needs systematic efforts to ensure their effectiveness and reproducibility. Attacking the most cutting-edge APIs like the latest GPT-4 will be our future work.
Appendix H Efficiency of LLM-based Mutation
The mutation mechanism is critical for genetic algorithms, as it diversifies the searching space for the algorithm and enables effective optimization. The LLM-based mutation proposed in our method, whose computation efficiency relied on the response speed of the LLMs, was not conducted frequently in our method. As claimed in Appendix D, the mutation probability is $0.01$, which means that every sample in the algorithm only has a 1% probability to have an LLM-base mutation.
We also show the computational cost in the ablation study (Table 5), which shows that the algorithm is even 5% faster when conducting LLM-based mutation ($722.6$s v.s. $766.6$s). This is because, as we mentioned above, the LLM-base mutation diversifies the searching space for the algorithm, making the algorithm achieve the success prompt earlier to trigger the termination criteria in Sec. 3.5. As for the effectiveness of attack, the LLM-based mutation introduces about 35% improvement as shown in Table 5. We believe the LLM-based mutation does introduce improvement for our method. However, we notice that there may exist a better mutation policy that can be further explored in future works.
For the effectiveness of the diverse initialization, we conducted an additional experiment with 100 data pairs from AdvBench Behaviour. In this experiment, we use AutoDAN-GA as the baseline and replace the LLM-based diverse initialization with a synonym replacement according to Wordnet. The synonym replacement is conducted in a probability=10% for every word in the prototype prompt, randomly picking a word from its synonyms. The results on Llama2 are as follows:
Table 10: Attack performance on different mutation schemes
Metric AutoDAN-GA w LLM-based diverse initializations AutoDAN-GA w wordnet synonym replacement
Keyword ASR 0.51 0.33
The results show that using a LLM to diversify the prompts surpasses naive synonym replacement. We take further case studies and find that the LLM-based diversification is often more reasonable and fluent. We believe this benefits from the outstanding power of the in-context understanding of the LLMs.
Appendix I Additional Defenses
We implement additional defenses that are proposed in Jain et al. (2023), including paraphrasing, and adversarial training. These defenses are implemented based on the paper’s description and safeguard a Vicuna-7b model. For paraphrasing, we use OpenAI’s GPT-3.5-turbo-0301 as the paraphraser. For adversarial training, we set mixing=0.2, epochs=3. We test the attack effectiveness of GCG attack and AutoDAN-HGA on 100 data pairs from the AdvBench Behavior dataset. Results are as follows:
Table 11: Attack performance against paraphrasing and adversarial training in Jain et al. (2023)
Keyword ASR Vicuna Vicuna+paraphrasing Vicuna+adversarial training
GCG 0.97 0.06 0.94
AutoDAN-HGA 0.97 0.68 0.93
A major difference is the performance of both attacks when facing the paraphrasing defense, where GCG becomes invalid, but AutoDAN retains its effectiveness on many samples. We take a further investigation and find that the paraphrasing usually “discards” the GCG’s suffixes as they are usually garbled characters. However, the jailbreak prompts generated by AutoDAN maintain their meaning after paraphrasing, as they inherently are meaningful texts.